10 Liters of a 20% acid solution are mixed with 30 Liters of a 30% acid solution. What is the concentration of acid in the final mixture?
Remember:
Concentration = (the number of Liters of pure acid) / (total number of Liters in the mixture)

Answer by stanbon(75887)

10 Liters of a 20% acid solution are mixed with 30 Liters of a 30% acid solution. What is the concentration of acid in the final mixture?
----
Equation:
acid + acid = acid
0.20*10 + 0.30*30 = x(40)
---
Multiply thru by 100 to get:
20*10 + 30*30 = 4000x
200 + 900 = 4000x
x = 1100/4000
x = 0.275 or 27.5%
=======================
Cheers,
Stan H.
=============